Algebraically solve: 70=6+8(^3x-1)
Answer by jainenderkapoor(61)

Algebraically solve: 70=6+8(^3x-1)
70 - 6 = 8^(3x -1)
64 = 8^(3x -1)
8^2 = 8^(3x -1)
Comparing both sides we get -- exponents to be same
2 = 3x - 1
2 + 1 = 3x
3 = 3x
x = 3/3 = 1
Hope the solution is clear to you. If you need any help, I am available for Online Tutoring at very reasonable rates. My email is is jainenderkapoor@gmail.com